Company use absorption costs approach for its sales and cost analysis

what is break even point/revenue? No beginning inventories so the production equals the sales volume, Variable and fixed cost are all manufacturing costs

BUDGET Apple Strawberry
Selling price 5.00 5.50
Variable cost 2.70 2.90
Fixed cost                       9,000                 4,500
Sales Volume                     25,000               10,000
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Calculation of break even point/revenue:
BUDGET Apple Strawberry Note
Selling price             5.00             5.50 A
Variable cost             2.70             2.90 B
Contribution per unit             2.30             2.60 C=A-B
Fixed cost      9,000.00     4,500.00 D
Break even point (units)      3,913.04     1,730.77 E=D/C
Break even point ($) 19,565.22     9,519.23 F=E*A